HETAV PATEL

Grade 11,

Calculate the molar mass of the following substances. Hydrochloric acid, HCl

 Calculate the molar mass of the following substances.
 Hydrochloric acid, HCl

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Hydrochloric acid, HCl = 1 × 1 + 1 × 35.5
= 1 + 35.5 = 36.5 g

Thanks